# Stress out of charmonia

### v1

@@ VI Summary and outlook
From these two primary quantities, we also define physical densities on the light front, including the pressure $\mathcal{P}(r_{\perp})$, the (proper) energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ and trace scalar density $\theta(r_{\perp})$. We also identify the positive-defined density $\mathcal{A}(r_{\perp})$, the Fourier transform of GFF $A(Q^{2})$, as the (convective) matter density. The energy density $\mathcal{E}(r_{\perp})$ is also positive and is consistent with the weak energy condition. These densities provide rich information of the system. For example, we find that there is a hierarchy for the sizes of the system: $r_{A}<r_{E}<r_{M^{2}}<r_{\theta}$, implying an onion-like structure of hadrons. We also show that there is a tachyonic core inside hadrons. Finally, we investigate the virial scaling index $2\langle T\rangle/\langle V\rangle$, which provides a quantitative measure of the strong force within charmonium. The obtained values are consistent with the non-relativistic picture.
@@
One of the interacting questions is the comparison between GFF $D$ extracted from $T^{+-}$ and from $T^{12}$ and from $T^{11}+T^{22}$. Operators $T^{11},T^{22}$ also contain interactions and need to be renormalized. However, for practical calculations on the light front, including perturbative or non-perturbative, counterterms computed from the light-front Hamiltonian are not sufficient to renormalize these operators. This is in sharp contrast to $T^{+-}$, which shares the same counterterms as its conserved charge, the light-front Hamiltonian. $T^{12}$ is interaction free. It is tempting to extract GFF $D$ from $T^{12}$ [55]. However, covariant light-front dynamics analysis shows that $T^{11},T^{22},T^{12}$ are all associated with a spurious GFF – a Lorentz structure that breaks the Poincaré symmetry. The GFF $D$ extracted from $T^{12}$ is likely to differ from what we obtained from $T^{+-}$, which does not suffer from the contamination of the spurious GFFs.
@@ Acknowledgements
The authors acknowledge fruitful discussions with C. Mondal, S. Nair, K. Serafin and J. More.

### v2

@@ VI Summary and outlook
From these two primary quantities, we also define physical densities on the light front, including the pressure $\mathcal{P}(r_{\perp})$, the (proper) energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density (aka. internal light-front energy density) $\mathcal{M}^{2}(r_{\perp})$ and trace scalar density $\theta(r_{\perp})$. We also identify the positive-defined density $\mathcal{A}(r_{\perp})$, the Fourier transform of GFF $A(Q^{2})$, as the (convective) matter density. The energy density $\mathcal{E}(r_{\perp})$ is also positive and is consistent with the weak energy condition. These densities provide rich information of the system. For example, we find that there is a hierarchy for the sizes of the system: $r_{A}<r_{E}<r_{M^{2}}<r_{\theta}$, implying an onion-like structure of hadrons. Finally, we investigate the virial scaling index $2\langle T\rangle/\langle V\rangle$, which provides a quantitative measure of the strong force within charmonium. The obtained values are consistent with the non-relativistic picture.
@@
One of the interesting questions is the comparison between GFF $D$ extracted from $T^{+-}$ and from $T^{12}$ and from $T^{11}+T^{22}$. Operators $T^{11},T^{22}$ also contain interactions and need to be renormalized. However, for practical calculations on the light front, including perturbative or non-perturbative, counterterms computed from the light-front Hamiltonian are not sufficient to renormalize these operators. This is in sharp contrast to $T^{+-}$, which shares the same counterterms as its conserved charge, the light-front Hamiltonian. $T^{12}$ is interaction free. It is tempting to extract GFF $D$ from $T^{12}$ [55]. However, covariant light-front dynamics analysis shows that $T^{11},T^{22},T^{12}$ are all associated with a spurious GFF – a structure that breaks the Poincaré symmetry. The GFF $D$ extracted from $T^{12}$ is likely to differ from what we obtained from $T^{+-}$, which does not suffer from the contamination of the spurious GFFs.
@@ Acknowledgements
The authors acknowledge fruitful discussions with C. Mondal, S. Nair, K. Serafin, J. More and C. Lorcé.